# CostFilter-AD: Enhancing Anomaly Detection through Matching Cost Filtering

### v1

@@ 1 Introduction
Unsupervised anomaly detection (UAD) plays a critical role in industrial quality inspection (Wu et al., 2024) by identifying anomalies at both image- and pixel-level using models trained solely on normal samples (Zhao, 2023; Chen et al., 2025). Being able to handle the scarcity and diversity of anomalies, as well as address the long-tail distribution of rare anomaly types through anomaly synthesis (Zhao, 2022) have been favored in particular. Existing approaches predominantly follow a “single model per category” paradigm (Liu et al., 2023; Zhang et al., 2023a), while effective, they often result in higher training costs and reduced scalability as anomaly categories expand. Thus, multi-class UAD with a unified model has emerged as a more scalable approach (Lu et al., 2023; He et al., 2024b; Yao et al., 2024). However, the inherent challenges of this task persist due to the diverse characteristics of anomalies across categories, particularly for subtle anomalies around small areas, with low contrast, or proximity to normals, making the unified detection extremely challenging.
@@
Inspired by the concept of matching cost filtering (also known as cost volume filtering) from the fields like stereo matching (Hosni et al., 2012; Kendall et al., 2017), depth estimation (Kam et al., 2022), and flow estimation (Gudovskiy et al., 2022), we reformulate anomaly detection as a three-step paradigm: feature extraction, anomaly cost volume construction, and anomaly cost volume filtering. With this idea, we propose CostFilter-AD, a generic post-processing plug-in for enhancing the anomaly detection of both reconstruction-based and embedding-based methods. Conceptually, we introduce a matching cost volume to address “what to match” and a cost volume filtering network to address “how to refine,” enabling adaptive noise suppression and more accurate matching between the input image and its corresponding templates.
@@
Our contributions are as follows: (i) We rethink UAD via matching cost volume filtering to explicitly tackle the intrinsic matching noise – an overlooked yet critical element with existing UAD methods. Under this perspective, we reformulate UAD with a three-step pipeline: feature extraction, matching cost volume construction, and cost volume filtering. (ii) We propose a novel method, CostFilter-AD, characterized by employing multi-layer input observations as attention queries to guide match denoising while preserving edge structures of subtle anomalies. Serving as a general plug-in, it can be seamlessly integrated with both reconstruction-based and embedding-based methods. (iii) Extensive experiments show that our method achieves new state-of-the-art performance under both multi-class and single-class UAD settings on the MVTec-AD and VisA benchmarks.
@@ 2.1 Unsupervised Anomaly Detection
UAD methods are broadly categorized as embedding-, reconstruction-, and synthesis-based methods (Cao et al., 2024). Embedding-based methods utilize pre-trained models for feature extraction, leveraging techniques like teacher-student networks (Deng & Li, 2022), distribution modeling (Defard et al., 2021; Li et al., 2023), or memory banks (Roth et al., 2022), but often struggle with adaptability to rare anomalies due to reliance on datasets like ImageNet (Deng et al., 2009). Reconstruction-based methods, including autoencoders (Gong et al., 2019), GANs (Liang et al., 2023; Lv et al., 2024), transformers (You et al., 2022), and diffusions (Zhang et al., 2023a), aim to rebuild normal patterns, but frequently contend with the “identical shortcut” issue. Synthesis-based methods generate pixel- or feature-level pseudo-anomalies (Zhao, 2022; Chen et al., 2025) to approximate real-world distributions but remain constrained by domain gaps (Zhang et al., 2025; Wang et al., 2023). Additionally, discriminative networks (Zavrtanik et al., 2021; Zhao, 2023) compare image pairs to detect anomalies, yet matching noise remains unresolved.
@@ 3.4 Anomaly Detection Output Generation
As shown in Fig. 2 (d), following stereo matching (Wang et al., 2024), the filtered anomaly volume undergoes global min pooling along the matching dimension, followed by a convolutional layer and softmax to generate the normal-anomaly score map for localization, denoted as $\mathcal{M}=\text{softmax}(\text{conv}(\text{min}(x)))$. As for detection, the image-level score is the average of the top $250$ values of the anomaly score map $\mathcal{M}$.
@@ 3.5 Training and Inference
The matching cost filtering process is designed as a normal-abnormal segmentation task, where the generated anomaly score maps $\mathcal{M}$ are supposed to align with the synthesized anomaly masks $\mathcal{M}_{s}$. Additionally, the class-aware adaptor outputs classification logits $\hat{Y}_{c}$ to enhance multi-class segmentation. The overall objective is defined as:
@@
where the parameter $\gamma$ in $\mathcal{L}_{\text{Focal}}$ is adjusted as $\gamma=\gamma_{0}-\sigma(\hat{Y}_{c})$ when the adaptor classifies correctly, and $\gamma=\gamma_{0}$ otherwise. Here, $\mathcal{L}_{\text{Focal}}$ is the focal loss to addresses the normal-anomaly imbalance, $\mathcal{L}_{\text{Soft-Iou}}$ refines anomaly region localization, $\mathcal{L}_{\text{SSIM}}$ preserves structural consistency, and $\mathcal{L}_{\text{CE}}$ enhances multi-class classification.
The inference process similarly constructs and filters the matching cost volume, yielding $\mathcal{M}$ for pixel-level localization and image-level score for detection.
@@ 4 Experimental Evaluation
Datasets. (1) MVTec-AD (Bergmann et al., 2019) is a widely used benchmark dataset for evaluating UAD models in industrial applications. It comprises 5,324 high-resolution images across 10 objects and 5 texture categories. Each category provides anomaly-free training samples and test images with various surface defects. (2) VisA (Zou et al., 2022) is a large-scale anomaly detection dataset with 10,821 high-resolution images spanning 12 subsets, including 9,621 normal and 1,200 abnormal images with surface defects (e.g., dents, scratches, color spots, and cracks) and structural anomalies such as misalignment or missing components.
Evaluation metrics. For a comprehensive evaluation, we adopt widely used metrics following prior works. For anomaly detection, we report image-level AUROC (I-AUROC), Average Precision (I-AP), and F1max (I-F1max). For anomaly localization, we employ pixel-level AUROC (P-AUROC), Average Precision (P-AP), F1max (P-F1max), and Area Under the Per-Region-Overlap Curve (AU-PRO). In the main text, we primarily report image/pixel AUROC, while other metrics are detailed in the appendix.
Implementation details. To validate our method as a plug-in solution, we integrate it with three recent multi-class UAD methods: GLAD (Yao et al., 2024) (diffusion-based), HVQ-Trans (Lu et al., 2023) (transformer-based) and AnomalDF (Damm et al., 2025) (embedding-based) under the full-shot setting, with full access to the training data set. We adhere to the original settings for a fair comparison. (i) Input images are resized to $256\times 256$ for GLAD and AnomalDF, and $224\times 224$ for HVQ-Trans. (ii) For the first two baselines, three templates ($N=3$) are randomly selected either from 25 diffusion denoising steps or the same-category training set. For HVQ-Trans, $N=1$ as it does not reconstruct intermediates. (iii) Multi-layer features ($L=4$) are extracted using a pre-trained DINO (Caron et al., 2021) with ViT-B/8 for GLAD and AnomalDF, and a pre-trained EfficientNet-B4 (Tan & Le, 2021) for HVQ-Trans. (iv) To reduce memory usage, the anomaly cost volume is trimmed to the top $D$ smallest channels as input for the first two baselines, aligning with HVQ-Trans, where the volume inherently has $D$ channels with a single template. (v) We train our models from scratch for 40 epochs with a batch size of 8, using the Adam optimizer initialized with a learning rate of $1\times 10^{-3}$. (vi) The loss weight $\alpha$ is set to 0.1 by default.
@@
Multiple Instances Macaroni1 90.5 / 93.3 92.6 / 98.6 85.7 / 94.1 97.5 / 99.6 91.6 / 99.5 93.1 / 99.9 96.0 / 99.9 88.7 / 99.1 93.7 / 99.4 88.0 / 98.2 95.3 / 99.9
@@
Single Instance Cashew 82.5 / 94.1 88.6 / 95.0 91.5 / 90.9 90.0 / 98.0 94.5 / 94.3 92.6 / 86.2 95.7 / 93.5 92.2 / 98.7 94.1 / 99.3 95.1 / 99.2 96.0 / 99.6
@@ 4.1 Quantitative Comparison
We reproduced the results of GLAD, HVQ-Trans, and AnomalDF, and integrated our method into these multi-class UAD baselines to validate its effectiveness. Moreover, we conducted a comprehensive comparison with various advanced methods, including synthetic-based JNLD (Zhao, 2022), CNN-based OmniAL (Zhao, 2023), diffusion-based DiAD (He et al., 2024b) and VPDM (Li et al., 2024), and Mamba-based MambaAD (He et al., 2024a).
@@
Multi-class UAD on MVTec-AD. As shown in Table 1, integrating our method yields significant AUROC improvements of 1.2%/0.9% for GLAD, 1.0%/0.7% for HVQ-Trans, and 1.7%/0.7% for AnomalDF in image- and pixel-level detection, achieving state-of-the-art results. More metrics in Table 7 and Table 8 further validate its effectiveness. Notably, in texture anomalies like the grid category, our method surpasses baselines by 4.9%/1.4%, 2.3%/1.3%, and 1.8%/1.7%, respectively. We attribute this to the method’s effective mitigation of matching noise, a critical yet often overlooked challenge that broadly exists in reconstruction- and embedding-based methods. These results highlight its strong generalization and superior performance across diverse anomaly classes.
@@
Table 4: Ablation studies of Glad+Ours on MVTec-AD. “$DN$$\rightarrow$ depth/channel” refers to mapping the matching dimension into the depth/channel dimension of the 3D U-Net. $\mathcal{C}_{0}$ denotes the volume uisng the final denoising step, $\mathcal{C}_{N-1}$ indicates uisng $N-1$ intermediate steps. SG and MG denote dual-stream attention guidance. $\mathcal{L}_{\text{F}}$ is focal loss, $\mathcal{L}_{\text{CE}}$ corresponds to the class-aware adaptor, and $\mathcal{L}_{\text{S}}$ is the combination of $\mathcal{L}_{\text{SSIM}}$ and $\mathcal{L}_{\text{Soft-Iou}}$.
$DN$$\rightarrow$ $DN$$\rightarrow$ channel $\mathcal{L}_{\text{F}}$ $\mathcal{L}_{\text{CE}}$ $\mathcal{L}_{\text{S}}$ Results
@@ 4.3 Ablation Studies on Each Component and Losses
To evaluate the contributions of each component and loss term in our method, we conducted ablation studies on the MVTec-AD benchmark using the Glad+Ours setting. Table 4 highlights the following key findings: (i) When directly adhering to stereo matching methods designed for local pixel-level matching, results drop to 87.8%/89.0%, likely due to feature contamination caused by global matching across multiple templates. In contrast, mapping the matching correspondence into the channel dimension boosts detection performance effectively. (ii) Using $\mathcal{C}_{0}$, which relies on the final denoised image as a template, achieves 96.2%/96.8%, while incorporating $N-1$ randomly selected intermediate denoised images further improves performance by 0.5%/0.5%. (iii) The proposed dual-stream attention guidance mechanism significantly enhances the filtering network’s learning. SG boosts spatial anomaly attention to 97.8%/97.5%, while the initial anomaly map-guided attention (MG) further improves channel matching, achieving 98.3%/97.8%. (iv) Loss functions: Focal loss serves as a fundamental loss, while $\mathcal{L}_{\text{CE}}$ improves performance to 98.5%/98.0%. Structural similarity and soft-IoU losses enhance structural consistency, and joint optimization yields a peak performance of 98.7%/98.2%. Overall, these findings underscore the complementary roles of each component and loss function in enhancing anomaly detection and localization, ensuring robustness and precision.
@@
97.5$\downarrow$ / 97.1$\downarrow$ 93.2 / 98.1
92.6$\downarrow$ / 98.0$\downarrow$
Embed. 94.5$\downarrow$ / 98.0$\downarrow$ 98.5
@@
85.6$\downarrow$ / 96.9$\downarrow$ 94.3 / 99.2
Hybrid 98.8$\uparrow$ / 98.1
98.6$\uparrow$ / 98.9$\uparrow$ 93.1
/ 98.2$\uparrow$

### v4

@@ 1 Introduction
Unsupervised anomaly detection (UAD) plays a critical role in industrial quality inspection (Wu et al., 2024) by identifying anomalies at both image- and pixel-level using models trained solely on normal samples (Zhao, 2023; Chen et al., 2025). Being able to handle the scarcity and diversity of anomalies, as well as address the long-tail distribution of rare anomaly types through anomaly synthesis (Zhao, 2022) have been favored in particular. Existing approaches predominantly follow a “single model per category” paradigm (Liu et al., 2023; Zhang et al., 2023; Chen et al., 2022), while effective, they often result in higher training costs and reduced scalability as anomaly categories expand. Thus, multi-class UAD with a unified model has emerged as a more scalable approach (Lu et al., 2023; He et al., 2024b; Yao et al., 2024). However, the inherent challenges of this task persist due to the diverse characteristics of anomalies across categories, particularly for subtle anomalies around small areas, with low contrast, or proximity to normals, making the unified detection extremely challenging.
@@
Inspired by the concept of matching cost filtering (also known as cost volume filtering) from the fields like stereo matching (Hosni et al., 2012; Kendall et al., 2017), depth estimation (Kam et al., 2022), flow estimation (Gudovskiy et al., 2022), and light field rendering (Wu et al., 2025), we reformulate anomaly detection as a three-step paradigm: feature extraction, anomaly cost volume construction, and anomaly cost volume filtering. With this idea, we propose CostFilter-AD, a generic post-processing plug-in for enhancing the anomaly detection of both reconstruction-based and embedding-based methods. Conceptually, we introduce a matching cost volume to address “what to match” and a cost volume filtering network to address “how to refine,” enabling adaptive noise suppression and more accurate matching between the input image and its corresponding templates.
@@
Our contributions are as follows: (i) We rethink UAD via matching cost volume filtering to explicitly tackle the intrinsic matching noise – an overlooked yet critical element with existing UAD methods. Under this perspective, we reformulate UAD with a three-step pipeline: feature extraction, matching cost volume construction, and cost volume filtering. (ii) We propose a novel method, CostFilter-AD, characterized by employing multi-layer input observations as attention queries to guide match denoising while preserving edge structures of subtle anomalies. Serving as a general plug-in, it can be seamlessly integrated with both reconstruction-based and embedding-based methods. (iii) Extensive experiments show that our method achieves new state-of-the-art performance under both multi-class and single-class UAD settings on multiple benchmarks.
@@ 2.1 Unsupervised Anomaly Detection
UAD methods are broadly categorized as embedding-, reconstruction-, and synthesis-based methods (Cao et al., 2024). Embedding-based methods utilize pre-trained models for feature extraction, leveraging techniques like teacher-student networks (Deng & Li, 2022), distribution modeling (Defard et al., 2021; Li et al., 2023; Lee & Choi, 2024), or memory banks (Roth et al., 2022; Bae et al., 2023), but often struggle with adaptability to rare anomalies due to reliance on datasets like ImageNet (Deng et al., 2009). Reconstruction-based methods, including autoencoders (Gong et al., 2019), GANs (Liang et al., 2023; Lv et al., 2024), transformers (You et al., 2022), diffusions (Zhang et al., 2023), and MoE (Meng et al., 2024), aim to rebuild normal patterns, but frequently contend with the “identical shortcut” issue. Synthesis-based methods generate pixel- or feature-level pseudo-anomalies (Zhao, 2022; Chen et al., 2025) to approximate real-world distributions but remain constrained by domain gaps (Zhang et al., 2025; Wang et al., 2023). Additionally, discriminative networks (Zavrtanik et al., 2021; Zhao, 2023) compare image pairs to detect anomalies, yet matching noise remains unresolved.
@@ 3.4 Anomaly Detection Output Generation
As shown in Fig. 2 (d), following stereo matching (Wang et al., 2024), the filtered anomaly volume undergoes global min pooling along the matching dimension, followed by a convolutional layer and softmax to generate the normal-anomaly score map for localization, denoted as $\mathcal{M}=\text{softmax}(\text{conv}(\text{min}(x)))$. As for detection, the image-level score is the average of the top $250$ values of the anomaly score map.
@@ 3.5 Training and Inference
The matching cost filtering process is designed as a normal-abnormal segmentation task, where the generated anomaly score maps $\mathcal{M}$ are supposed to align with the synthesized anomaly masks $\mathcal{M}_{s}$. The training objective is defined as:
@@
where $\mathcal{L}_{\text{Focal}}$ is the focal loss to addresses the normal-anomaly imbalance, $\mathcal{L}_{\text{Soft-Iou}}$ refines anomaly region localization, $\mathcal{L}_{\text{SSIM}}$ preserves structural consistency, and $\mathcal{L}_{\text{CE}}$ enhances multi-class classification. Notably, the parameter $\gamma$ in $\mathcal{L}_{\text{Focal}}$ is adjusted as $\gamma=\gamma_{0}-\sigma(\hat{Y}_{c})$ when the adaptor classifies correctly, and $\gamma=\gamma_{0}$ otherwise. Thus, the class-aware adaptor modulates $\gamma$ via its predicted logits $\hat{Y}_{c}$ to enhance multi-class segmentation.
The inference process similarly constructs and filters the matching cost volume, yielding anomaly map $\mathcal{M}$. To integrate with the baseline response, we compute a weighted sum $\lambda\cdot\mathcal{M}+(1-\lambda)\cdot\mathcal{M}_{\text{baseline}}$ for anomaly localization and detection, where $\lambda\in[0,1]$ compensates for potential scale differences between the two components.
@@ 4 Experimental Evaluation
Datasets. (1) MVTec-AD (Bergmann et al., 2019) is a widely used dataset for evaluating UAD models in industrial applications. It comprises 5,324 high-resolution images across 10 objects and 5 texture categories. Each category provides anomaly-free training samples and test images with various surface defects. (2) VisA (Zou et al., 2022) is a large-scale dataset comprising 10,821 high-resolution images spanning 12 subsets, including 9,621 normal and 1,200 anomalous images. It covers a wide range of surface defects (e.g., dents, scratches, color spots, and cracks) as well as structural anomalies such as misalignment and missing components. (3) MPDD (Jezek et al., 2021) is a challenging dataset consisting of 1,346 images from 6 metal part categories. It includes 888 normal training samples and 458 test images, with pixel-level annotations for anomaly regions. (4) BTAD (Mishra et al., 2021) comprises 2,830 images from three industrial categories for UAD.
Evaluation metrics. For a comprehensive evaluation, we adopt widely used metrics following prior works. For anomaly detection, we report image-level AUROC (I-AUROC), AUPRC (I-AP), and F1max (I-F1max). For anomaly localization, we employ pixel-level AUROC (P-AUROC), AUPRC (P-AP), F1max (P-F1max), and Area Under the Per-Region-Overlap Curve (AUPRO). In the main text, we primarily report image/pixel AUROC, while other metrics are detailed in the appendix.
Implementation details. To validate our method as a plug-in solution, we integrate it with three recent multi-class UAD methods: GLAD (Yao et al., 2024) (diffusion-based), HVQ-Trans (Lu et al., 2023) (transformer-based) and AnomalDF (Damm et al., 2025) (embedding-based) under the full-shot setting, with full access to the training data set. We adhere to the original settings for a fair comparison. (i) Input images are resized to $256\times 256$ for GLAD and AnomalDF, and $224\times 224$ for HVQ-Trans. (ii) For the first two baselines, three templates ($N=3$) are randomly selected either from 25 diffusion denoising steps or the same-category training set. For HVQ-Trans, $N=1$ as it does not reconstruct intermediates. (iii) In our setting, AnomalDF (+Ours) refers to a variant of AnomalyDINO that dynamically samples templates from the full training set for each input, in contrast to the original fixed few-shot protocol with a static template pool, thereby achieving a trade-off between template diversity and memory efficiency. (iv) Multi-layer features ($L=4$) are extracted using a pre-trained DINO (Caron et al., 2021) with ViT-B/8 for GLAD and AnomalDF, and a pre-trained EfficientNet-B4 (Tan & Le, 2021) for HVQ-Trans. (v) To reduce memory usage, the anomaly cost volume is trimmed to the top $D$ smallest channels as input for the first two baselines, aligning with HVQ-Trans, where the volume inherently has $D$ channels with a single template. (vi) We train our models from scratch for 40 epochs with a batch size of 8, using the Adam optimizer initialized with a learning rate of $1\times 10^{-3}$. (vii) The loss weight $\alpha$ is set to 0.1 by default.
@@
Multiple Instances
Macaroni1 90.5 / 93.3 92.6 / 98.6 85.7 / 94.1 97.5 / 99.6 91.6 / 99.5 93.1 / 99.9 96.0 / 99.9 88.7 / 99.1 93.7 / 99.4 88.0 / 98.2 95.3 / 99.9
@@
Single Instance
Cashew 82.5 / 94.1 88.6 / 95.0 91.5 / 90.9 90.0 / 98.0 94.5 / 94.3 92.6 / 86.2 95.7 / 93.5 92.2 / 98.7 94.1 / 99.3 95.1 / 99.2 96.0 / 99.6
@@
Table 3: Multi-class UAD evaluation across additional baselines and benchmarks using seven metrics, reporting category-wise mean results for each benchmark.
Benchmark Method Image-level Pixel-level
AU-ROC AP F1max AU-ROC AP F1max AUPRO
MVTec-AD UniAD 97.5 99.1 97.0 96.9 44.5 50.5 90.6
+Ours 99.0 99.7 98.1 97.5 60.5 59.9 91.3
Dinomaly 99.6 99.8 99.0 98.3 68.7 68.7 94.6
+Ours 99.7 99.8 99.1 98.4 68.9 68.9 94.8
VisA UniAD 91.5 93.6 88.5 98.0 32.7 38.4 76.1
+Ours 92.1 94.0 88.9 98.6 34.0 39.0 86.4
Dinomaly 98.7 98.9 96.1 98.7 52.5 55.4 94.5
+Ours 98.7 99.0 96.3 98.8 53.2 55.8 94.7
MPDD HVQ-Trans 86.5 87.9 85.6 96.9 26.4 30.5 88.0
+Ours 93.1 95.4 90.3 97.5 34.1 37.0 82.9
Dinomaly 97.3 98.5 95.6 99.1 60.0 59.8 96.7
+Ours 97.5 98.5 95.8 99.2 60.2 59.9 96.7
BTAD HVQ-Trans 90.9 97.8 94.8 96.7 43.2 48.7 75.6
+Ours 93.3 98.6 96.0 97.3 47.0 50.2 76.2
Dinomaly 95.4 98.5 95.5 97.9 70.1 68.0 76.5
+Ours 95.5 98.6 95.8 98.1 74.3 69.8 77.5
@@ 4.1 Quantitative Comparison
We reproduced the results of UniAD (You et al., 2022), GLAD, HVQ-Trans, AnomalDF, and Dinomaly (Guo et al., 2025), and integrated our method into these multi-class UAD baselines to validate its effectiveness. Moreover, we conducted a comprehensive comparison with various advanced methods, including synthetic-based JNLD (Zhao, 2022), CNN-based OmniAL (Zhao, 2023), diffusion-based DiAD (He et al., 2024b) and VPDM (Li et al., 2024), and Mamba-based MambaAD (He et al., 2024a).
Multi-class UAD on MVTec-AD. As shown in Table 1, integrating our method yields significant AUROC improvements of 1.2%/0.9% for GLAD, 1.0%/0.7% for HVQ-Trans, and 1.7%/0.7% for AnomalDF in image- and pixel-level detection, achieving state-of-the-art results. More metrics in Table 9 to Table 12 further validate its effectiveness. Notably, in texture anomalies like the grid category, our method surpasses baselines by 4.9%/1.4%, 2.3%/1.3%, and 1.8%/1.7%, respectively. We attribute this to the method’s effective mitigation of matching noise, a critical yet often overlooked challenge that broadly exists in reconstruction- and embedding-based methods. These results highlight its strong generalization and superior performance across diverse anomaly classes.
@@
Multi-class UAD across more baselines and benchmarks. To further assess the generalizability of our plug-in design, we evaluate CostFilter-AD on a broader set of recent multi-class UAD baselines and benchmarks, using seven comprehensive metrics at both image and pixel levels. As summarized in Table 3, we present category-wise mean results for each benchmark to provide an overall performance perspective; detailed per-class metrics are available in Table 9 to Table 20. Our method consistently improves performance across all baselines, datasets, and evaluation dimensions. For example, when integrated with Dinomaly (Guo et al., 2025), CostFilter-AD achieves AUROC scores of 99.6%/98.5% (image/pixel) on MVTec-AD, 98.8%/98.8% on VisA, 97.4%/99.1% on MPDD, and 95.8%/98.1% on BTAD. We attribute these consistent gains to its ability to effectively filter matching noise that is widespread yet often overlooked, while preserving subtle anomaly structures and features across a wide range of anomaly categories.
@@
Table 5: Ablation studies of Glad+Ours on MVTec-AD. $DN\rightarrow$ depth/channel” refers to mapping the matching dimension into the depth/channel dimension of the 3D U-Net. $\mathcal{C}_{0}$ denotes the volume uisng the final denoising step, $\mathcal{C}_{N-1}$ indicates uisng N1N-1 intermediate steps. SG and MG denote dual-stream attention guidance. $\mathcal{L}_{\text{F}}$ is focal loss, $\mathcal{L}_{\text{CE}}$ corresponds to the class-aware adaptor, and $\mathcal{L}_{\text{S}}$ is the combination of $\mathcal{L}_{\text{SSIM}}$ and $\mathcal{L}_{\text{Soft-Iou}}$.
$\rightarrow DNDN\rightarrow$ channel $\mathcal{L}_{\text{F}}$ $\mathcal{L}_{\text{CE}}$ $\mathcal{L}_{\text{S}}$ Results
@@ 4.3 Ablation Studies on Each Component and Losses
To evaluate the contributions of each component and loss term in our method, we conducted ablation studies on the MVTec-AD benchmark using the Glad+Ours setting. Table 5 highlights the following key findings: (i) When directly adhering to stereo matching methods designed for local pixel-level matching, results drop to 87.8%/89.0%, likely due to feature contamination caused by global matching across multiple templates. In contrast, mapping the matching correspondence into the channel dimension boosts detection performance effectively. (ii) Using $\mathcal{C}_{0}$, which relies on the final denoised image as a template, achieves 96.2%/96.8%, while incorporating N1N-1 randomly selected intermediate denoised images further improves performance by 0.5%/0.5%. (iii) The proposed dual-stream attention guidance mechanism significantly enhances the filtering network’s learning. SG boosts spatial anomaly attention to 97.8%/97.5%, while the initial anomaly map-guided attention (MG) further improves channel matching, achieving 98.3%/97.8%. (iv) Loss functions: Focal loss serves as a fundamental loss, while $\mathcal{L}_{\text{CE}}$ improves performance to 98.5%/98.0%. Structural similarity and soft-IoU losses enhance structural consistency, and joint optimization yields a peak performance of 98.7%/98.2%. Overall, these findings underscore the complementary roles of each component and loss function in enhancing anomaly detection and localization, ensuring robustness and precision.
@@
[…]
@@
[…]
Hybrid $8\uparrow / 9$8.1
[…]